378 𝑙𝑝𝑠 of air is isothermally compressed at 27℃ and 1379 π‘˜π‘ƒπ‘Žπ‘Ž; 𝑝2 = 4136 π‘˜π‘ƒπ‘Žπ‘Ž. For both nonflow and steady-flow (𝑉1 = 23 π‘š 𝑠 , 𝑉2 = 46 π‘š 𝑠 , βˆ†π‘ƒ = 0) processes, compute (a) ∫ 𝑝𝑑𝑉 π‘Žπ‘›π‘‘ βˆ’ ∫ 𝑉𝑑𝑝 (b) βˆ†π‘ˆ, βˆ†π», π‘Žπ‘›π‘‘ βˆ†π‘† (c) π‘Š π‘Žπ‘›π‘‘ 𝑄 


1
Expert's answer
2021-11-01T12:54:50-0400

a)

"\int pdV=\nu RT=572.5~kJ,"

"\int Vdp=p_2V_2-p_2V_2=-575.2~kJ,"

b)

"\Delta U=p\Delta V-V\Delta p=4.8~kJ,"

"\Delta H=\Delta U+pV=572.5~kJ,"

"\Delta S=\frac UT=1.9085~\frac{kJ}K,"

c)

"W=2\Delta U=9.6~kJ,"

"Q=pV=575.2~kJ."